1) A Class contains 10 men and 20 women of which half of the men and half of the women have brown eyes. Find the probability that a person chosen at random is a man or has brown eyes.
Answer by stanbon(75887)

A Class contains 10 men and 20 women of which half of the men and half of the women have brown eyes. Find the probability that a person chosen at random is a man or has brown eyes.
--------
P(A or B) = P(A) + P(B) - P(A and B)
--------------
P(man or brown eyes) = P(man) + P(brown eyes) - P(man and brown eyes)
P(m or be) = (10/20) + (10/20) - (5/30)
= 1- (5/30) = 25/30 = 5/6
==============================
Cheers,
Stan H.